OpenStudy (kyleighpaige):

how do you factor 2x^2 + 4x + 18

Elsa213 (elsa213):

\(\LARGE\color{blue}{2x ^2 +4x+18}\)

Elsa213 (elsa213):

That's the problem, right?

OpenStudy (kyleighpaige):

Yes that's the problem.

OpenStudy (jiteshmeghwal9):

Split the middle term such that the product of its components is equal to 36x^2 and the sum of the components is equal to 4x

Elsa213 (elsa213):

>.>

OpenStudy (mathmale):

I'd advise that you reduce the expression by dividing all terms by 2. Your result?

OpenStudy (mathmale):

Unfortunately, there's ambiguity in your presentation of this problem. If you write "2x squared," others may interpret that as (1) (2x)^2 or as (2) 2*x^2. Which do you mean? Leave no doubt regarding your intentions.

OpenStudy (bthompson361):

Use the quadratic formula \[-b + or - \sqrt{b ^{2}-4(ac)}/2a\]

OpenStudy (bgarland2003):

Are you allowed to use the quadratic formula or do you have to factor

OpenStudy (mathmale):

As given to you: 2x^2 + 4x + 18 After factoring out the '2:' x^2 + 2x + 9 Focus on factoring x^2 + 2x + 9. Suggest using the quadratic formula. If you do, you will come up with two complex roots (such as "a + i*b"). If, strictly for example, you come up with the root 2-ib, then a factor would be x-2+ib. Focus on factoring x^2 + 2x + 9. Suggest using the quadratic formula.

OpenStudy (mathmale):

Quadratic formula, correct version: \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (mathmale):

Not quite correct, because of ambiguity in regard to what is being divided by 2a: \[-b + or - \sqrt{b ^{2}-4(ac)}/2a\]

OpenStudy (naka354):

2*x^2 + 4*x + 18 = 2*(x^2 + 2*x + 9)